40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6     Entry time: Sat Oct 20 11:54:13 2007
Author: waldman 
Type: Other 
Category: OMC 
Subject: OMC and OMC-SUS work 
[Rich, Chub, Pinkesh, Chris, Sam]

Friday the 18th was a busy day in OMC land. Both DCPDs were mounted to the glass breadboard and the OMC-SUS structure was rebuilt to the point that an aluminum dummy mass is hanging, unbalanced. The OSEMs have not be put on the table cloth yet, but everything is hanging free. As for the DCPDs, if you recall one beam is 3mm off center from the DCPD tombstone. Fortunately, one DCPD is nearly 3mm offcenter from the case in the right direction, so the errors nearly cancel. The DCPD is too high, so the beam isn't quite centered, but they're close. We'll get photos of the beam positions in someday. Also, the DC gain between the two PDs is, at first glance, different by 15%. DCPD1, the one seen in transmission has 315 mV of signal while DCPD2 has 280 mV. Not sure why, could be because of beam alignment or tolerances in the Preamp or the angle incident on the diode or the QE of the diodes. The glass cans have *not* been removed.
ELOG V3.1.3-